# Approximation schemes for covering and packing mixed-integer programs with a fixed number of constraints

### v1

@@ Abstract
This paper presents an algorithmic study of a class of covering mixed-integer linear programming problems which encompasses classic cover problems, including multidimensional knapsack, facility location and supplier selection problems. We first show some properties of the vertices of the associated polytope, which are then used to decompose the problem into instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. The proposed decomposition is used to design a polynomial-time approximation scheme for the problem with a fixed number of constraints. To the best of our knowledge, this is the first approximation scheme for such a general class of covering mixed-integer programs. Moreover, we design a fully polynomial-time approximation scheme and an approximate linear programming formulation for the case with a single constraint. These results improve upon the previously best-known 2-approximation algorithm for the knapsack cover problem with a single continuous variable. Finally, we show a perfect compact formulation for the case where all variables have the same lower and upper bounds. Analogous results are derived for the packing and assignment variants of the problem.
@@ 1 Introduction
This paper presents an algorithmic study of mixed-integer programming problems, denoted by p, which are defined as:
@@
$y_{i}\in\{0,1\}$ $\forall i\in[n].$ (4)
where $m$ and $n$ are positive integers, $v,\ell,c\in\mathbb{Z}_{\geq}^{nm}$, $d\in\mathbb{Z}_{\geq}^{m}$ and $f\in\mathbb{Z}_{\geq}^{n}$ denote the nonnegative integer parameters and $a]=\{1,\ldots,a\}$ for a positive integer $a$. Throughout this paper, it is assumed, without loss of generality, that $d_{j}\geq c_{ij}\geq\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$. For completeness, we also consider the packing variant of p, defined analogously as:
@@
The covering integer programs (cip) are closely related to p. Given a matrix $A\in\mathbb{R}^{m\times n}_{\geq}$, vectors $a\in\mathbb{R}^{m}_{\geq}$, and $b,h\in\mathbb{R}^{n}_{\geq}$, cip is defined as $\min\{h^{T}z\mathrel{\mathop{\ordinarycolon}}Az\geq a,z\leq b,z\in\mathbb{Z}^{n}_{\geq}\}$. carr2000strengthening propose a $\Delta_{1}$-approximation for this problem, where $\Delta_{1}$ is the maximum number of non-zero coefficients in any row of $A$. kolliopoulos2005approximation design a $\mathcal{O}(\log m)$-approximation for cip, and observe that this is asymptotically the best possible unless $\mathsf{P}=\mathsf{NP}$. This inapproximability threshold follows from the classic set cover problem, which does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$ (raz1997sub). The set cover problem is a special case of cip where $z$ is binary, and binary cip is a particular case of p where $\ell_{ij}=c_{ij}=A_{ij}$, $v_{ij}=0$, and $f_{i}=h_{i}$ for all $i\in[n]$ and $j\in[m]$. As a consequence, p does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$. Although there is little hope of designing better approximations for p in general, the case where the number of constraints is fixed remains largely unexplored. Observe that p with fixed $m$ encompasses classic cover problems, including multidimensional knapsack, facility location and supplier selection problems. Therefore, p remains $\mathsf{NP}$-hard even when $m=1$. For any fixed $m\geq 2$p with $m$ constraints does not admit a fully polynomial-time approximation scheme (FPTAS) unless $\mathsf{P}=\mathsf{NP}$. This is due to the analogous hardness result for multidimensional knapsack cover that can be obtained following essentially the same proof for the multidimensional knapsack (packing) devised by magazine1984note, and to the fact that multidimensional knapsack cover is a particular case of p with a fixed number of constraints. On the positive side, friezeApproximationAlgorithmsMdimensional1984 design a polynomial-time approximation scheme (PTAS) to the multidimensional knapsack cover problem. For the knapsack cover problem (i.e., the one-dimensional case), guntzer2000approximate propose an FPTAS which is based on a greedy strategy.
@@
The packing integer programs (pip) are defined analogously and encompass classic combinatorial optimization problems such as knapsack, matching, and independent set problems. As a consequence, binary pip (and thus the packing variant of p) cannot be approximated within $n^{1-\epsilon}$, for any $\epsilon>0$, unless $\mathsf{P}=\mathsf{NP}$ (zuckerman2006linear). There exists a vast literature on approximation algorithms for binary pip, nearly all of which rely on randomized rounding techniques (see, e.g., srinivasan1995improved; bansal2012solving; chekuri2004multidimensional).
[…]
@@
For p with a single constraint, we devise an FPTAS based on a dynamic programming approach in Section 4. A byproduct of this result is an FPTAS for the knapsack cover with a single continuous variable, thereby improving upon the previously best-known 2-approximation algorithm by zhaoApproximationAlgorithms012014. We note that vanhoeselFullyPolynomialApproximation2001 present an FPTAS for the single-item capacitated economic lot-sizing problem that is more general than one-dimensional p in its consideration of backlogging and holding costs. However, their model does not consider lower bounds. Additionally, for each $\epsilon\in(0,1)$, we prove the existence of a compact linear programming formulation which yields a solution of value at most $(1+\epsilon)$ times the optimal value of the problem.
@@
Finally, in Section 5, we show a perfect compact formulation for the one-dimensional p with uniform bounds, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. constantinoLowerBoundsLotSizing1998a investigates the uncapacitated version of this problem, i.e., the case where $c\geq d$. More precisely, he derives two families of valid inequalities and shows that they describe the convex hull in this special case. Although one-dimensional uniform p is a special case (with positive demand only in the last period) of a single-item lot-sizing problem with piecewise linear costs, which is known to be polynomial-time solvable (hellionPolynomialTimeAlgorithm2012a), no tight formulation is known for this problem. Hence the proposed perfect compact formulation contributes in this direction. Concluding remarks and directions for further investigation are presented in Section 6. Note that each section also includes an analysis of the corresponding packing problem.
@@ 2 Decomposition into multidimensional knapsack problems
For the converse, define a vector $(x,y)\in\mathbb{R}^{nm}\times\{0,1\}^{n}$ that equals zero except for the entries $y_{i}=1$ for all $i\in[n]$ with $f_{i}=0$, and $x_{ij}=c_{ij}$ for all $j\in[m]$ and $i\in Z_{j}$. If follows from the construction that $(x,y)$ is a feasible solution of $I$ of value equal to zero. Hence, $\operatorname{OPT}(I)=0$. ∎
@@
We next present properties of an optimal solution of p. These allow us to demonstrate how to decompose p into a collection of instances of the following generalization of the multidimensional knapsack cover problem.
@@
Output: Subset $S\subseteq[\eta]$, and nonnegative $\alpha_{j}\leq\bar{c}_{j}$ for all $j\in[\mu]$ such that $\sum_{i\in S}\bar{w}_{ij}\geq d_{j}-\alpha_{j}$ for every $j\in[\mu]$.
@@
Note that an instance of mkc is feasible if and only if $\sum_{i\in[\eta]}\bar{w}_{ij}\geq d_{j}-\bar{c}_{j}$ for every $j\in[\mu]$. We assume henceforth that all considered instances of mkc are feasible.
@@
Lemma 2.2.
There exists an optimal solution $(x,y)$ to p such that the following assertions hold for every $j\in[m]$.
  1. (i)
    There exists at most one $g(j)\in[n]$ such that $x_{g(j),j}\in(\ell_{g(j),j},c_{g(j),j})$. Precisely, $\ell_{g(j),j}<x_{g(j),j}<c_{g(j),j}$ if any exists. Otherwise, let $g(j)\in\arg\max_{i\in[n]}\{v_{ij}\mathrel{\mathop{\ordinarycolon}}x_{ij}=c_{ij},y_{i}=1\}$ if $x_{ij}\in\{\ell_{ij}y_{i},c_{ij}y_{i}\}$ for all $i\in[n]$ and there is at least one $i\in[n]$ with $x_{ij}=c_{ij}$, and $g(j)\in\arg\min_{i\in[n]}\{v_{ij}\mathrel{\mathop{\ordinarycolon}}x_{ij}=\ell_{ij},y_{i}=1\}$ otherwise. We define $g(j)$ as the fractional variable for each $j\in[m]$.
  2. (ii)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$, and $x_{ij}=c_{ij}$ otherwise.
  3. (iii)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}=v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $i<g(j)$, and $x_{ij}=c_{ij}$ otherwise.
Proof.
Let $(x,y)\in\mathbb{R}^{nm+n}$ be an optimal solution to p that is an extreme point of $conv(X)$, and let $j\in[m]$. Suppose to the contrary that $\ell_{aj}<x_{aj}<c_{aj}$ and $\ell_{bj}<x_{bj}<c_{bj}$ for some $a,b\in[n]$ with $a\neq b$. Let $\epsilon=\min_{d\in\{a,b\}}\min\{c_{dj}-x_{dj},\>x_{dj}-\ell_{dj}\}$. We define two vectors $x^{1},x^{2}\in\mathbb{R}^{nm+n}$ such that $x^{1}$ and $x^{2}$ are equal to $x$ except for the following entries: $x_{aj}^{1}=x_{aj}+\epsilon$, $x_{bj}^{1}=x_{bj}-\epsilon$, $x_{aj}^{2}=x_{aj}-\epsilon$, and $x_{bj}^{2}=x_{bj}+\epsilon$. It is clear that $(x^{1},y)$ and $(x^{2},y)$ belong to $X$. Moreover, $(x,y)=\frac{1}{2}(x^{1},y)+\frac{1}{2}(x^{2},y)$, a contradiction to the choice of $(x,y)$. Hence, there exists at most one $a\in[n]$ with $\ell_{aj}<x_{aj}<c_{aj}$.
Suppose first that such $a$ exists and let $g(j)=a$, i.e., $\ell_{g(j),j}<x_{g(j),j}<c_{g(j),j}$. Suppose to the contrary that $y_{i}=1$, $x_{ij}=c_{ij}>\ell_{ij}$ and $v_{ij}>v_{g(j),j}$ for some $i\in[n]\setminus\{g(j)\}$. Define a vector $\bar{x}\in\mathbb{R}^{nm}$ to be equal to $x$ except for $\bar{x}_{ij}=x_{ij}-\epsilon$ and $\bar{x}_{g(j),j}=x_{g(j),j}+\epsilon$, where $\epsilon\mathrel{\mathop{\ordinarycolon}}=\min\{c_{g(j),j}-x_{g(j),j},\>x_{ij}-\ell_{ij}\}$. It follows that $(\bar{x},y)$ is a feasible solution to p of cost strictly smaller than $(x,y)$, a contradiction. Analogously, one can prove that $x_{ij}=\ell_{ij}$ if $y_{i}=1$ and $v_{ij}>v_{g(j),j}$. Suppose now that $x_{ij}\in\{\ell_{ij}y_{i},c_{ij}y_{i}\}$ for all $i\in[n]$. Note that $v_{ij}\geq v_{i^{\prime}j}$ for all $i,i^{\prime}\in[n]$ such that $y_{i}=y_{i^{\prime}}=1$, $x_{ij}=\ell_{ij}<c_{ij}$ and $x_{i^{\prime}j}=c_{i^{\prime}j}>\ell_{i^{\prime}j}$. Otherwise, one could obtain a solution cheaper than $(x,y)$ by increasing $x_{ij}$ and decreasing $x_{i^{\prime}j}$ by the same (sufficiently small) amount. Let $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$. If $x_{g(j),j}=c_{g(j),j}$, then the previous observation together with the definition of $g(j)$ in (i) imply that $x_{ij}=c_{ij}$ if $v_{ij}<v_{g(j),j}$, and $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$. If $x_{g(j),j}=\ell_{g(j),j}$, then there is no $i\in[n]$ with $y_{i}=1$ and $x_{ij}=c_{ij}$. Thus (ii) follows from the choice of $g(j)$.
Consider now $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}=v_{g(j),j}$. We define the set of inversions of any solution $(x^{\prime},y^{\prime})$ satisfying (i) as
Let $V$ be the set of inversion of $(x,y)$. Suppose that $(x,y)$ has at least one inversion (i.e., $V\neq\emptyset$). Assume, without loss of generality, that this inversion is due to an element in $[n]$ smaller than $g(j)$, and choose $i\in V$ to be the largest number such that $i<g(j)$. For each $i^{\prime}\in\{i+1,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, starting from $g(j)$ to $i+1$, we first move from $x_{ij}$ to $x_{i^{\prime}j}$ the maximum quantity $\epsilon$ such that either $x_{ij}-\epsilon=\ell_{ij}$ or $x_{i^{\prime}j}+\epsilon=c_{i^{\prime}j}$. Then we are done in the former case, or we repeat the procedure with $i^{\prime}-1$ or end the procedure when $i^{\prime}-1=i$ in the latter one. Let $\bar{x}\in\mathbb{R}^{nm}$ be the vector obtained at the end of this procedure. Note that there is $t\in\{i,i+1,\ldots,g(j)\}$ with $y_{t}=1$ and $v_{tj}=v_{g(j),j}$ such that, for each $i^{\prime}\in\{i,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, we have $\bar{x}_{i^{\prime}j}=\ell_{i^{\prime}j}$ if $i^{\prime}<t$ and $\bar{x}_{i^{\prime}j}=c_{i^{\prime}j}$ if $i^{\prime}>t$. Hence, the number of inversions of $(\bar{x},y)$ is strictly smaller than the number of inversions of $(x,y)$. Moreover, since $v_{i^{\prime}j}=v_{g(j),j}$ for all $i^{\prime}\in V$, it is clear that $(\bar{x},y)$ is an optimal solution to p that still respects (i) and (ii). By repeating the previous procedure at most n1n-1 times, we obtain an optimal solution to p which satisfies (i), (ii), and (iii). The case of an inversion involving an element in $[n]$ larger than $g(j)$ is handled analogously.
Let $j\in[m]$ and $k\in[n]$. We define $L_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}>v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i<k\}$ and $C_{kj}=\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}<v_{kj}\}\cup\{i\in[n]\setminus\{k\}\mathrel{\mathop{\ordinarycolon}}v_{ij}=v_{kj}\text{ and }i>k\}.$ Note that $\{L_{kj},C_{kj}\}$ is a partition of $n]\setminus\{k\}$ which only depends on the input instance, and can be easily computed in linear time.
@@
The proof is immediate from Lemma 2.2 and the definitions of the $L$ and $C$ sets. ∎
@@
$w_{ij}^{g}=\begin{cases}\ell_{ij}&\text{ if $i\in L_{g(j),j}$,}\\ c_{ij}&\text{ if $i\in C_{g(j),j}$,}\\ \ell_{g(j)j}&\text{ if $i=g$.}\end{cases}$
@@
$\min\;\;$ $\sum_{g\in G}\left(\sum_{i\in[n]}f^{g}_{i}y_{i}^{g}+\sum_{j\in[m]}v_{j}^{g}\alpha_{j}^{g}\right)$ (5)
s.t. $\sum_{g\in G}z^{g}=1,$ (6)

### v2

# Covering and packing mixed-integer linear programs with a fixed number of constraints: Approximation and convex hull

@@ Abstract
This paper presents an algorithmic study of a class of covering mixed-integer linear programming problems which encompasses classic cover problems, including multidimensional knapsack, facility location and supplier selection problems. We first show some properties of optimal solutions, which are then used to decompose the problem into instances of the multidimensional knapsack cover problem with a single continuous variable per dimension. The proposed decomposition is used to design a polynomial-time approximation scheme for the problem with a fixed number of constraints. To the best of our knowledge, this is the first approximation scheme for such a general class of covering mixed-integer linear programs. Moreover, we design a fully polynomial-time approximation scheme and an approximate linear programming formulation for the case with a single constraint. These results improve upon the previously best-known 2-approximation algorithm for the knapsack cover problem with a single continuous variable. Finally, we show a perfect compact formulation for the case where all variables have the same lower and upper bounds. Analogous results are derived for the packing and more general variants of the problem.
@@ 1 Introduction
This paper presents an algorithmic study of mixed-integer linear programming problems, denoted by p, which are defined as:
@@
$y_{i}\in\{0,1\}$ $\forall i\in[n],$ (4)
where $m$ and $n$ are positive integers, $v,\ell,c\in\mathbb{Z}_{\geq}^{nm}$, $d\in\mathbb{Z}_{>}^{m}$ and $f\in\mathbb{Z}_{\geq}^{n}$ denote the nonnegative integer parameters with $c_{ij}\geq\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$ and $a]=\{1,\ldots,a\}$ for any positive integer $a$. This model is applicable to a range of problems, including variants of the knapsack problem and the supplier selection problem. In the latter, the aim is to cover the demand for $m$ items, where item $j\in[m]$ has demand $d_{j}$, by choosing a subset of suppliers from a given set of $n$ candidates, subject to minimum and maximum order quantities $\ell_{ij}$ and $c_{ij}$ for each supplier $i\in[n]$ and each item $j\in[m]$, and to minimize the total cost, where $f_{i}$ is the fixed cost of selecting supplier $i\in[n]$ and $v_{ij}$ is the unit cost of an item $j\in[m]$ ordered from supplier $i\in[n]$.
For completeness, we also consider the packing variant of p, defined analogously as:
@@
The covering integer programs (cip) are closely related to p. Given a matrix $A\in\mathbb{R}^{m\times n}_{\geq}$, vectors $a\in\mathbb{R}^{m}_{\geq}$, and $b,h\in\mathbb{R}^{n}_{\geq}$, cip is defined as $\min\{h^{T}z\mathrel{\mathop{\ordinarycolon}}Az\geq a,z\leq b,z\in\mathbb{Z}^{n}_{\geq}\}$. Carr et al. Carr et al. (2000) propose a $\Delta_{1}$-approximation for this problem, where $\Delta_{1}$ is the maximum number of non-zero coefficients in any row of $A$. Kolliopoulos and Young Kolliopoulos and Young (2005) design a $\mathcal{O}(\log m)$-approximation for cip, and observe that this is asymptotically the best possible unless $\mathsf{P}=\mathsf{NP}$. This inapproximability threshold follows from the classic set cover problem, which does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$ (Raz and Safra, 1997). The set cover problem is a special case of cip where $z$ is binary, and binary cip is a particular case of p where $\ell_{ij}=c_{ij}=A_{ij}$, $v_{ij}=0$, and $f_{i}=h_{i}$ for all $i\in[n]$ and $j\in[m]$. As a consequence, p does not admit a $o(\log m)$-approximation unless $\mathsf{P}=\mathsf{NP}$. Furthermore, for arbitrary $m>1$, no constant approximation ratio can be obtained for p unless $\mathsf{P}=\mathsf{NP}$ (Srinivasan, 1995).
Although there is little hope of designing better approximations for p in general, the case where the number of constraints is fixed remains largely unexplored. Observe that p with fixed $m$ encompasses classic cover problems, including multidimensional knapsack cover problems and therefore, p remains $\mathsf{NP}$-hard even when $m=1$. For any fixed $m\geq 2$p with $m$ constraints does not admit a fully polynomial-time approximation scheme (FPTAS) unless $\mathsf{P}=\mathsf{NP}$. This is due to the analogous hardness result for multidimensional knapsack cover that can be obtained following essentially the same proof for the multidimensional knapsack (packing) devised by Magazine and Chern Magazine and Chern (1984), and to the fact that multidimensional knapsack cover is a particular case of p with a fixed number of constraints. On the positive side, Frieze and Clarke Frieze and Clarke (1984) design a polynomial-time approximation scheme (PTAS) to the multidimensional knapsack cover problem. Kulik and Shachnai Kulik and Shachnai (2010) show that no efficient PTAS (EPTAS) exists for this problem unless $W[1]=$ FPT. For the knapsack cover problem (i.e., the one-dimensional case), Güntzer and Jungnickel Güntzer and Jungnickel (2000) propose an FPTAS which is based on a greedy strategy.
@@
The packing integer programs (pip) are defined analogously and encompass classic combinatorial optimization problems such as knapsack, matching, and independent set problems. As a consequence, binary pip (and thus the packing variant of p) cannot be approximated within $n^{1-\epsilon}$, for any $\epsilon>0$, unless $\mathsf{P}=\mathsf{NP}$ (Zuckerman, 2006). There exists a vast literature on approximation algorithms for binary pip, nearly all of which rely on randomized rounding techniques (see, e.g., Bansal et al., 2012; Chekuri and Khanna, 2004; Srinivasan, 1995).
We also investigate a more general class of packing mixed-integer programs, defined as follows:
s.t. $\sum_{i\in[n]}x_{ij}\leq d_{j}$ $\forall j\in[m],$
$\sum_{j\in[m]}y_{ij}\leq t_{i}$ $\forall i\in[n],$
$\ell_{ij}y_{ij}\leq x_{ij}\leq c_{ij}y_{ij}$ $\forall i\in[n],\;j\in[m],$
$y_{ij}\in\{0,1\}$ $\forall i\in[n],\;j\in[m],$
where $t\in\mathbb{Z}^{n}_{>}$ and the other parameters are defined as before. Note that this includes relevant problems as the Generalized Assignment Problem (gap). In particular, when $t_{i}=1$, $v_{ij}=0$ and $c_{ij}=\ell_{ij}$ for all $i\in[n]$ and $j\in[m]$, this problem is equivalent to gap, which is known to be $\mathsf{APX}$-hard (Chekuri and Khanna, 2005). On the positive side, for any $\epsilon>0$, there is an approximation for gap with ratio $(e/(e-1)+\epsilon)\approx(1.582+\epsilon)$ using an LP-rounding approach, which is currently the best known approximation for this problem (Fleischer et al., 2006).
Contributions
This work provides a theoretical study of the three problem classes introduced above through the lens of approximation algorithms and perfect linear programming formulations. In Section 2, we first show some properties of some optimal solutions of p. Using these properties, we decompose p into instances of a generalization of the multidimensional knapsack cover problem with one continuous variable per dimension. The proposed decomposition of p is used to design a PTAS for p with a fixed number of constraints in Section 3. To the best of our knowledge, this is the first approximation scheme for such a general class of covering mixed-integer linear programs. Furthermore, in a sense, it is the best possible approximation for this problem since p admits no FPTAS, as previously observed. In Section 4, using a similar strategy, we also design PTASes for the packing variants of p.
@@
For p with a single constraint, we devise an FPTAS based on a dynamic programming approach in Section 5. A byproduct of this result is an FPTAS for the knapsack cover with a single continuous variable, thereby improving upon the previously best-known 2-approximation algorithm by Zhao and Li Zhao and Li (2014). We note that Van Hoesel and Wagelmans Van Hoesel and Wagelmans (2001) present an FPTAS for the single-item capacitated economic lot-sizing problem that is more general than one-dimensional p in its consideration of backlogging and holding costs. However, their model does not consider lower bounds. Additionally, for each $\epsilon\in(0,1)$, we prove the existence of a compact linear programming formulation that yields a solution of value at most $(1+\epsilon)$ times the optimal value of the problem.
@@
Finally, in Section 6, we provide a perfect compact formulation for the one-dimensional p with uniform bounds, that is, $\ell_{i}=\ell$ and $c_{i}=c$ for all $i\in[n]$. Constantino Constantino (1998) investigates the uncapacitated version of this problem, i.e., the case where $c\geq d$. More precisely, he derives two families of valid inequalities and shows that they describe the convex hull in this special case. Although one-dimensional uniform p is a special case (with positive demand only in the last period) of a single-item lot-sizing problem with piecewise linear costs, which is known to be polynomial-time solvable (Hellion et al., 2012), no tight formulation is known for this problem. Hence the proposed perfect compact formulation contributes in this direction. Concluding remarks and directions for further investigation are presented in Section 7. Note that each section also includes an analysis of the corresponding packing problem.
@@ 2 Decomposition into multidimensional knapsack problems
For the converse, define a vector $(x,y)\in\mathbb{R}^{nm}\times\{0,1\}^{n}$ that equals zero except for the entries $y_{i}=1$ for all $i\in[n]$ with $f_{i}=0$, and $x_{ij}=c_{ij}$ for all $j\in[m]$ and $i\in Z_{j}$. If follows from the construction that $(x,y)$ is a feasible solution of $I$ of value equal to zero. Hence, $\operatorname{OPT}(I)=0$.
@@
We next present properties of an optimal solution of p, which allow us to decompose p into instances of the following generalization of the multidimensional knapsack cover problem. In this problem, we have $\eta$ items to choose from to cover $\mu$ dimensions, where item $i\in[\eta]$ has weight $\bar{w}_{ij}$ in dimension $j\in[\mu]$, the cost of choosing item $i\in[\eta]$ is $\bar{f}_{i}$ and the demand to be covered in dimension $j\in[\mu]$ is $\bar{d}_{j}$. In addition, we can use at most $\bar{c}_{j}$ amount of a resource to cover the demand in dimension $j\in[\mu]$ at a cost of $\bar{v}_{j}$ per unit. The aim is to cover the demand in all dimensions at minimum cost. More precisely,
@@
Output: Subset $S\subseteq[\eta]$, and nonnegative $\alpha_{j}\leq\bar{c}_{j}$ for all $j\in[\mu]$ such that $\sum_{i\in S}\bar{w}_{ij}\geq\bar{d}_{j}-\alpha_{j}$ for every $j\in[\mu]$.
@@
Note that an instance of mkc is feasible if and only if $\sum_{i\in[\eta]}\bar{w}_{ij}\geq\bar{d}_{j}-\bar{c}_{j}$ for every $j\in[\mu]$. We assume henceforth that all considered instances of mkc are feasible.
@@
Let $(x,y)\in\mathbb{R}^{nm+n}$ be an optimal solution to p that is an extreme point of $conv(X)$, and let $j\in[m]$. Define $N_{1}=\{i\in[n]\mathrel{\mathop{\ordinarycolon}}y_{i}=1,x_{ij}=\ell_{ij}\}$, $N_{2}=\{i\in[n]\mathrel{\mathop{\ordinarycolon}}y_{i}=1,x_{ij}=c_{ij}\}$, and $N_{3}=\{i\in[n]\mathrel{\mathop{\ordinarycolon}}\ell_{ij}<x_{ij}<c_{ij}\}$. Note that we cannot have $N_{1}=N_{2}=N_{3}=\emptyset$ as $d>0$.
Suppose to the contrary that $\ell_{aj}<x_{aj}<c_{aj}$ and $\ell_{bj}<x_{bj}<c_{bj}$ for some $a,b\in[n]$ with $a\neq b$. Let $\epsilon=\min_{i\in\{a,b\}}\min\{c_{ij}-x_{ij},\>x_{ij}-\ell_{ij}\}$. We define two vectors $x^{1},x^{2}\in\mathbb{R}^{nm+n}$ such that $x^{1}$ and $x^{2}$ are equal to $x$ except for the following entries: $x_{aj}^{1}=x_{aj}+\epsilon$, $x_{bj}^{1}=x_{bj}-\epsilon$, $x_{aj}^{2}=x_{aj}-\epsilon$, and $x_{bj}^{2}=x_{bj}+\epsilon$. It is clear that $(x^{1},y)$ and $(x^{2},y)$ belong to $X$. Moreover, $(x,y)=\frac{1}{2}(x^{1},y)+\frac{1}{2}(x^{2},y)$, a contradiction to the choice of $(x,y)$. Hence, there exists at most one $a\in[n]$ with $\ell_{aj}<x_{aj}<c_{aj}$, and so $|N_{3}|\leq 1$.
We define $g(j)=i$ if $N_{3}=\{i\}$. Otherwise, it holds that $N_{3}=\emptyset$, and we define
$g(j)=\begin{cases}\displaystyle\arg\max_{i\in N_{2}}\{v_{ij}\}&\parbox[t]{238.49231pt}{\raggedright if $N_{2}\neq\emptyset$,\@add@raggedright}\\[10.0pt] \displaystyle\arg\min_{i\in N_{1}}\{v_{ij}\}&\text{otherwise}.\end{cases}$
We next prove that the following two assertions hold.
  1. (i)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$ and $x_{ij}=c_{ij}$ otherwise.
  2. (ii)
    For each $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}=v_{g(j),j}$, we have $x_{ij}=\ell_{ij}$ if $i<g(j)$ and $x_{ij}=c_{ij}$ otherwise.
Suppose first that $\ell_{g(j),j}<x_{g(j),j}<c_{g(j),j}$. Suppose to the contrary that $y_{i}=1$, $x_{ij}=c_{ij}>\ell_{ij}$ and $v_{ij}>v_{g(j),j}$ for some $i\in[n]\setminus\{g(j)\}$. Define a vector $\bar{x}\in\mathbb{R}^{nm}$ to be equal to $x$ except for $\bar{x}_{ij}=x_{ij}-\epsilon$ and $\bar{x}_{g(j),j}=x_{g(j),j}+\epsilon$, where $\epsilon\mathrel{\mathop{\ordinarycolon}}=\min\{c_{g(j),j}-x_{g(j),j},\>x_{ij}-\ell_{ij}\}$. It follows that $(\bar{x},y)$ is a feasible solution to p of cost strictly smaller than $(x,y)$, a contradiction. Analogously, one can prove that $x_{ij}=\ell_{ij}$ if $y_{i}=1$ and $v_{ij}>v_{g(j),j}$. Suppose now that $x_{ij}\in\{\ell_{ij}y_{i},c_{ij}y_{i}\}$ for all $i\in[n]$. Note that $v_{ij}\geq v_{i^{\prime}j}$ for all $i,i^{\prime}\in[n]$ such that $y_{i}=y_{i^{\prime}}=1$, $x_{ij}=\ell_{ij}<c_{ij}$ and $x_{i^{\prime}j}=c_{i^{\prime}j}>\ell_{i^{\prime}j}$. Otherwise, one could obtain a solution cheaper than $(x,y)$ by increasing $x_{ij}$ and decreasing $x_{i^{\prime}j}$ by the same (sufficiently small) amount. Let $i\in[n]\setminus\{g(j)\}$ such that $y_{i}=1$ and $v_{ij}\neq v_{g(j),j}$. If $x_{g(j),j}=c_{g(j),j}$, then the previous observation together with the definition of $g(j)$ imply that $x_{ij}=c_{ij}$ if $v_{ij}<v_{g(j),j}$, and $x_{ij}=\ell_{ij}$ if $v_{ij}>v_{g(j),j}$. If $x_{g(j),j}=\ell_{g(j),j}$, then there is no $i\in[n]$ with $y_{i}=1$ and $x_{ij}=c_{ij}$. Thus (i) follows from the choice of $g(j)$.
We define the set of inversions of any solution $(x^{\prime},y^{\prime})$ as
Let $V$ be the set of inversion of $(x,y)$. Suppose that $(x,y)$ has at least one inversion (i.e., $V\neq\emptyset$). Assume, without loss of generality, that this inversion is due to an element in $[n]$ smaller than $g(j)$, and choose $i\in V$ to be the largest number such that $i<g(j)$. For each $i^{\prime}\in\{i+1,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, starting from $g(j)$ to $i+1$, we first move from $x_{ij}$ to $x_{i^{\prime}j}$ the maximum quantity $\epsilon$ such that either $x_{ij}-\epsilon=\ell_{ij}$ or $x_{i^{\prime}j}+\epsilon=c_{i^{\prime}j}$. Then we are done in the former case, or we repeat the procedure with $i^{\prime}-1$ or end the procedure when $i^{\prime}-1=i$ in the latter one. Let $\bar{x}\in\mathbb{R}^{nm}$ be the vector obtained at the end of this procedure. Note that there is $t\in\{i,i+1,\ldots,g(j)\}$ with $y_{t}=1$ and $v_{tj}=v_{g(j),j}$ such that, for each $i^{\prime}\in\{i,\ldots,g(j)\}$ with $y_{i^{\prime}}=1$ and $v_{i^{\prime}j}=v_{g(j),j}$, we have $\bar{x}_{i^{\prime}j}=\ell_{i^{\prime}j}$ if $i^{\prime}<t$ and $\bar{x}_{i^{\prime}j}=c_{i^{\prime}j}$ if $i^{\prime}>t$. Hence, the number of inversions of $(\bar{x},y)$ is strictly smaller than the number of inversions of $(x,y)$. Moreover, since the procedure only changes entries $i\in[n]$ such that $v_{ij}=v_{g(j),j}$, it is clear that $(\bar{x},y)$ is an optimal solution to p that still respects (i). By repeating the previous procedure at most n1n-1 times, we obtain an optimal solution to p which satisfies (i), and (ii). The case of an inversion involving an element in $[n]$ larger than $g(j)$ is handled analogously.
@@
$w_{ij}^{g}=\begin{cases}\ell_{ij}&\text{ if $i\in L_{g(j),j}$,}\\ c_{ij}&\text{ if $i\in C_{g(j),j}$,}\\ \ell_{g(j)j}&\text{ if $i=g(j)$.}\end{cases}$
@@
$\min\;\;$ $\sum_{g\in\mathcal{G}}\left(\sum_{i\in[n]}f^{g}_{i}y_{i}^{g}+\sum_{j\in[m]}v_{j}^{g}\alpha_{j}^{g}\right)$ (5)
s.t. $\sum_{g\in\mathcal{G}}z^{g}=1,$ (6)